Problem 10-16 Comprehensive Variance Analysis Lo10-1, LO10-2, LO10-3] Highland Company produces a lightweight backpack that is popular with college students. Standard variable costs relating to a single backpack are given below: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials $6.00 per yard Direct labor Variable manufacturing overhead Total standard cost per unit per direct labor- $2 hour S ? Overhead is applied to production on the basis of direct labor-hours. During March, 520 backpacks were manufactured and sold Selected information relating to the month's production is given below Variable Materials Used 15,600 $ 14,325 Direct Manufacturing Labor $7,280 Overhead $1,456 $1,717 Total standard cost allowed* Actual costs incurred Materials price variance Materials quantity variance Labor rate variance Labor efficiency variance Variable overhead rate variance Variable overhead efficiency variance $1,590 U For the month's production. The following additional information is available for March's production: Actual direct labor-hours Difference between standard and actual cost per backpack 780 $ 0.20 F produced during March
What is the standard direct labor rate per hour? What was the labor rate variance for March? The labor efficiency variance? What was the variable overhead rate variance for March? The variable overhead efficiency variance? Prepare a standard cost card for one backpack.
Answer #1

Solution :

What is the standard direct labor rate per hour?

Particulars Amount
Standard variable manufacturing overhead cost for march $1456
Standard variable manufacturing overhead rate per direct labor hour $2
Standard direct labor hours for march

=1456/2

= $ 728

Standard direct labor rate per hour

=$ 7280/$728

= $ 10

What was the labor rate variance for march? The labor efficiency variance?

Particulars Amount Amount
Actual Cost per unit of back pack production

=(15,600+7280+1456)/(520)-(0.20)

$ 46.6
Total Number of produced Back packs 520
Total Actual cost of production $ 24,232
Less: Actual cost of materials 14,325
Actual cost of manufacturing Overhead 1717

=14325+1717

= $ 16042

Actual cost of Direct Labor

=24,232-16042

= $ 8,190

Direct labor Rate = (780*$10)-$ 8190 = -$ 390
Labor efficiency varaince = 7280-8190 = -$ 910

What was the variable overhead rate variance for march?The variable overhead efficiency variance?

Variable overhead rate variance=(Actual Hour*Actual Rate)-(Actual Hour of input*Standard Rate)

=(1717-(780*2))

= $ 157

Variable overhead rate variance = $ 157

Variable overhead efficiency variance ( Actual Hour * Standard rate)-(Standard Hour * Standard R ate)

=(780 * 2)-1456

=1560-1456

= $ 104

Prepare a standard cost card for one backpack?

Particulars SR SC of Back pack Amount
Direct Materials

=($15600/520)/$6

= 5 Yards

$ 6

= 5 * $ 6

= $ 30

Direct labor

= $ 728/520

= $ 1.4

$10 $14
Variable Manufacturing Overhead $ 1.4 $ 2

= $ 1.4 * $ 2

= $ 2.8

Total Standard cost per unit of Backpack

= $ 30+$ 14+$ 2.8

= $ 46.8